# Large-scale semi-organized rolls in a sheared convective turbulence: Mean-field simulations

### v1

@@ Abstract
Based on a mean-field theory of a non-rotating turbulent convection (Phys. Rev. E 66, 066305, 2002), we perform mean-field simulations (MFS) of sheared convection which takes into account an effect of modification of the turbulent heat flux by the non-uniform large-scale motions. This effect is caused by production of additional essentially anisotropic velocity fluctuations generated by tangling of the mean-velocity gradients by small-scale turbulent motions due to the influence of the inertial forces during the lifetime of turbulent eddies. These anisotropic velocity fluctuations contributes to the turbulent heat flux. As the result of this effect, there is an excitation of large-scale convective-shear instability, which causes the formation of large-scale semi-organized structures in the form of rolls. The life-times and spatial scales of these structures are much larger compared to the turbulent scales. By means of MFS performed for stress-free and no-slip vertical boundary conditions, we determine the spatial and temporal characteristics of these structures. Our study demonstrates that the modification of the turbulent heat flux by non-uniform flows leads to a strong reduction of the critical effective Rayleigh number (based on the eddy viscosity and turbulent temperature diffusivity) required for the formation of the large-scale rolls. During the nonlinear stage of the convective-shear instability, there is a transition from the two-layer vertical structure with two roles in the vertical direction before the system reaches steady-state to the one-layer vertical structure with one role after the system reaches steady-state. This effect is observed for all effective Rayleigh numbers. We find that inside the convective rolls, the spatial distribution of the mean potential temperature includes regions with a positive vertical gradient of the potential temperature caused by the mean heat flux of the convective rolls. This study might be useful for understanding of the origin of large-scale rolls observed in atmospheric convective boundary layers as well as in numerical simulations and laboratory experiments.
@@ I Introduction
Temperature stratified turbulence and turbulent convection exist in many geophysical and astrophysical flows as well as in industrial flows [see, e.g., Refs. 1, 2, 3, 4, 5, 6, 7]. In spite of turbulent transport has been studied more than 100 years, some key questions remain unclear due to extreme values of the governing parameters in geophysical and astrophysical flows [see, e.g., Refs. 8, 9, 10, 11, 12].
@@
  • The level of the mean kinetic energy at saturation increases with increase of the scale separation ratio, and it is very weekly dependent on the effective Rayleigh number.
@@ II Sheared turbulent convection and turbulent heat flux
$\biggl{(}{\partial\over\partial t}+\overline{\bm{U}}\cdot\mbox{\boldmath$\nabla$}\biggr{)}\overline{U}_{i}+(\overline{\bm{U}}_{\rm S}\cdot\mbox{\boldmath$\nabla$})\overline{U}_{i}+(\overline{\bm{U}}\cdot\mbox{\boldmath$\nabla$})\overline{\bm{U}}_{\rm S}$
$\quad\quad\quad=-{\nabla}_{i}\biggl{(}{\overline{P}\over\overline{\rho}_{0}}\biggr{)}-{\nabla}_{j}\langle u_{i}\,u_{j}\rangle+\beta\,\overline{\Theta}e_{i},$ (1)
$\biggl{(}{\partial\over\partial t}+\overline{\bm{U}}\cdot\mbox{\boldmath$\nabla$}\biggr{)}\overline{\Theta}+(\overline{\bm{U}}_{\rm S}\cdot\mbox{\boldmath$\nabla$})\overline{\Theta}=-(\overline{\bm{U}}\cdot\mbox{\boldmath$\nabla$})\overline{T}_{0}$
$\quad\quad\quad-{\nabla}_{i}\langle u_{i}\theta\rangle,$ (2)
where $\overline{\bm{U}}_{\rm S}=(S\,z,0,0)$ is the shear velocity directed along the $x$ axis, $\overline{\Theta}$ is the mean potential temperature defined as $\overline{\Theta}=\overline{T}\,(\overline{P}_{0}/\overline{P})^{1-1/\gamma}$. Here $\beta=|{\bm{g}}|/T_{0}$ is the buoyancy parameter, ${\bm{g}}=-g\,{\bm{e}}$ is the acceleration caused by the gravity, ${\bm{e}}$ is the unit vector in the vertical direction (along the $z$ axis), $\gamma=c_{\rm p}/c_{\rm v}$ is the specific heats ratio, $\overline{T}$ is the mean physical temperature with the reference value $\overline{T}_{0}$ as the temperature in the equilibrium (i.e., the basic reference state), $\overline{P}$ is the mean pressure with the reference value $\overline{P}_{0}$ and $\overline{\rho}_{0}$ is the mean fluid density in the equilibrium. For large Reynolds and Peclet numbers, we neglect in Eqs. (1)–(2) small terms due to the kinematic viscosity and molecular diffusivity of the potential temperature in comparison with those due to the turbulent viscosity and turbulent diffusivity. In Eqs. (1)–(2), the mean fields corresponds to deviations from the equilibrium: $\overline{\bm{U}}_{0}=\overline{\bm{U}}_{\rm S}$, $\;{\bm{\nabla}}\overline{P}_{0}=\overline{\rho}_{0}{\bm{g}}$ and $\overline{\rho}_{0}=$ const.
@@
where ${\bf F}_{x}^{\rm CW}=-\tau_{0}\,({\bf F}_{z}^{\ast}\cdot\mbox{\boldmath$\nabla$})\overline{\bm{U}}_{\rm S}(z)$ is the counter-wind turbulent heat flux and ${\bf F}^{\ast}=-\kappa_{{}_{T}}\mbox{\boldmath$\nabla$}\overline{\Theta}$ is the classical turbulent heat flux, $\tau_{0}$ is the correlation time of turbulent velocity at the integral scale of turbulent motions, $\overline{\bm{W}}=\mbox{\boldmath$\nabla$}{\bf\times}\overline{\bm{U}}$ is the mean vorticity, $\overline{\bm{U}}=\overline{\bm{U}}_{\perp}+\overline{\bm{U}}_{z}$ is the mean velocity with the horizontal $\overline{\bm{U}}_{\perp}$ and vertical $\overline{\bm{U}}_{z}$ components. The additional terms in the turbulent heat flux result in the excitation of large-scale instability and formation of the large-scale convective rolls [50, 51, 52].
@@
The physics related to the additional terms in the turbulent heat flux is discussed below. The term $\propto-\tau_{0}\,{\bm{F}}_{z}^{\ast}\,{\rm div}\,\overline{\bm{U}}_{\perp}$ in Eq. (3) for the turbulent heat flux causes the redistribution of the vertical background turbulent heat flux ${\bm{F}}_{z}^{\ast}$ by the perturbations of the convergent (or divergent) horizontal mean velocity $\overline{\bm{U}}_{\perp}$ (see Fig. 1) during the life-time of turbulent eddies. As the result, this additional contribution enhances the vertical turbulent flux of potential temperature due to the converging horizontal motions, which increases the buoyancy, thus creating the upward flow. The latter increases the horizontal convergent flow.
@@
On the other hand, the term $\propto(\tau_{0}/2)\,(\overline{\bm{W}}{\bm{\times}}{\bm{F}}_{z}^{\ast})$ in Eq. (3) produces the horizontal turbulent heat flux by the ”rotation” of the vertical background turbulent heat flux ${\bm{F}}_{z}^{\ast}$ caused by the perturbations of the horizontal mean vorticity $\overline{\bm{W}}_{\perp}$. This decreases local potential temperatures in rising motions, which decreases the buoyancy accelerations, and weakens vertical velocity and vorticity.
@@
The last term $\propto(\tau_{0}/2)\,\left(\overline{\bm{W}}_{z}{\bf\times}{\bf F}_{x}^{\rm CW}\right)$ in Eq. (3) for the turbulent heat flux produces the horizontal heat flux through the ”rotation” of the horizontal background counter-wind turbulent heat flux ${\bf F}_{x}^{\rm CW}$ by the vertical component of the mean vorticity. The counter-wind turbulent flux of potential temperature ${\bf F}_{x}^{\rm CW}$ arises due to the following reasons. In a horizontally homogeneous and sheared convective turbulence, the mean shear velocity $\overline{U}^{\rm S}_{x}(z)$ increases with the height, while the mean potential temperature $\overline{\Theta}(z)$ decreases with the height. Uprising fluid particles produce both, positive fluctuations of potential temperature, $\theta>0$ because $\partial\theta/\partial t\propto-({\bm{u}}\cdot\mbox{\boldmath$\nabla$})\overline{\Theta}(z)$, and negative fluctuations of horizontal velocity, $u_{x}<0$ because $\partial u_{x}/\partial t\propto-({\bm{u}}\cdot\mbox{\boldmath$\nabla$})\overline{U}^{\rm S}_{x}(z)$. This creates a negative horizontal turbulent flux of potential temperature, $u_{x}\,\theta<0$. On the other hand, sinking fluid particles create both, negative fluctuations of potential temperature, $\theta<0$, and positive fluctuations of horizontal velocity, $u_{x}>0$, resulting in negative horizontal turbulent flux of potential temperature, $u_{x}\,\theta<0$. Therefore, the net horizontal turbulent flux of potential temperature is negative, ${\bf F}_{x}^{\rm CW}\equiv\langle u_{x}\,\theta\rangle<0$, in spite of a zero horizontal mean temperature gradient. Therefore, the counter-wind turbulent flux of potential temperature modifies the turbulent potential temperature flux caused by the non-uniform mean velocity field. The counter-wind turbulent flux is associated with non-gradient turbulence transport of heat.
@@
The last term $\propto(\tau_{0}/2)\,\overline{\bm{W}}_{z}{\bf\times}{\bf F}_{x}^{\rm CW}$ in Eq. (3), causes generation of the cross-wind horizontal turbulent heat flux by turning the counter-wind horizontal turbulent flux ${\bf F}_{x}^{\rm CW}$ by perturbations of vertical component of the mean vorticity $\overline{\bm{W}}_{z}$ (see Fig. 2). This produces alternating pairs of convergence or divergence cross-wind turbulent heat fluxes, ${\bf F}_{y}^{\rm new}=(\tau_{0}/2)\,\overline{\bm{W}}_{z}{\bf\times}{\bf F}_{x}^{\rm CW}$, resulting alternative warmer or cooler patches which, in turn, cause alternative upward warm and downward cool motions. This is precisely the mechanism of large-scale instability responsible for formation of the large-scale convective rolls stretched along the mean shear and generation of convective-shear waves propagating perpendicular to the convective rolls in the sheared convection [50, 51, 52].
@@ III Governing equations and convective-shear instability
Refer to captionRefer to caption
Figure 3: The growth rate $\gamma_{\rm inst}$ of the instability (upper panel) and the frequency of the excited convective-shear waves $\omega$ (bottom panel) versus the parameter $\alpha$ for the effective Rayleigh number ${\rm Ra}_{{}_{T}}=0.5$ and different the non-dimensional shear number ${\rm Sh}=$ 0.03 (dashed); 0.1 (dashed-dotted); 0.3 (solid).
Refer to captionRefer to caption
Figure 4: The growth rate $\gamma_{\rm inst}$ of the instability (upper panel) and the frequency of the excited convective-shear waves $\omega$ (bottom panel) versus the parameter $\alpha$ for the effective Rayleigh number ${\rm Ra}_{{}_{T}}=10^{3}$ and different the non-dimensional shear number ${\rm Sh}=$ 0.03 (dashed); 0.1 (dashed-dotted); 0.3 (solid).
@@
${\partial\tilde{\bm{U}}\over\partial t}+(\tilde{\bm{U}}\cdot{\bm{\nabla}})\tilde{\bm{U}}+(\tilde{\bm{U}}_{\rm S}\cdot\mbox{\boldmath$\nabla$})\tilde{\bm{U}}+(\tilde{\bm{U}}\cdot\mbox{\boldmath$\nabla$})\tilde{\bm{U}}_{\rm S}$
$\;=-{{\bm{\nabla}}\tilde{P}\over\rho_{0}}+{\rm Ra}_{{}_{T}}\,\tilde{\Theta}\,{\bm{e}}+\Delta\tilde{\bm{U}},$ (4)
@@
where the non-dimensional mean velocity $\tilde{\bm{U}}$ with ${\rm div}\,\tilde{\bm{U}}=0$, the mean potential temperature $\tilde{\Theta}$ and the mean pressure $\tilde{P}$ are shown with tilde, the flux $\tilde{F}_{z}^{\ast}$ is the nondimensional vertical turbulent background heat flux, the unit vector ${\bm{e}}$ is directed along vertical $z$ axis, the non-dimensional shear velocity is $\tilde{\bm{U}}_{\rm S}=({\rm Sh}\,\tilde{z},0,0)$, and $\tilde{z}$ is is the non-dimensional vertical coordinate.
@@
where $S=\nabla_{z}\overline{U}^{\rm S}_{x}$ is the linear velocity shear, i.e., $S$ is constant, $u_{c}=(gF_{z}\ell_{0})^{1/3}$ is the convective velocity, $F_{z}$ is the vertical turbulent flux of potential temperature, $\kappa_{{}_{T}}$ is the turbulent (eddy) diffusivity and $\tau_{0}=\ell_{0}/u_{0}$. In Eqs (4)–(5) we have neglected small terms ${\rm O}(\epsilon^{2})$.
@@
Figure 7: Time evolution of the maximum velocity $\tilde{U}_{\rm max}(t)$ for the effective Rayleigh number ${\rm Ra}_{{}_{T}}=0.5$, $\alpha=0.5$ and shear number ${\rm Sh}=1$, and at different values of the parameter $\epsilon=10^{-3};5\times 10^{-3};10^{-2}$, for the stress-free boundary conditions.
Figure 8: Time evolution of the maximum velocity $\tilde{U}_{\rm max}(t)$ for the effective Rayleigh number ${\rm Ra}_{{}_{T}}=0.5$, $\alpha=0.5$ and shear number ${\rm Sh}=1$, and at different values of the parameter $\epsilon=10^{-3};5\times 10^{-3};10^{-2}$, for the no-slip boundary conditions.
Figure 9: Time evolution of the maximum velocity $\tilde{U}_{\rm max}(t)$ for the effective Rayleigh number ${\rm Ra}_{{}_{T}}=1000$, $\alpha=0.5$ and shear number ${\rm Sh}=1$, and at different values of the parameter $\epsilon=10^{-3};5\times 10^{-3};10^{-2}$, for the stress-free boundary conditions.
Figure 10: Time evolution of the maximum velocity $\tilde{U}_{\rm max}(t)$ for the effective Rayleigh number ${\rm Ra}_{{}_{T}}=1000$, $\alpha=0.5$ and shear number ${\rm Sh}=1$, and at different values of the parameter $\epsilon=10^{-3};5\times 10^{-3};10^{-2}$, for the no-slip boundary conditions.
@@
Numerical solution of Eq. (13) yields the growth rate $\gamma_{\rm inst}$ of the instability and the frequency $\omega$ of the excited convective-shear waves versus the parameter $\alpha$ for the effective Rayleigh numbers ${\rm Ra}_{{}_{T}}=0.5$ (see upper panel in Fig. 3) and ${\rm Ra}_{{}_{T}}=10^{3}$ (see upper panel in Fig. 4), and for different values of the non-dimensional shear number ${\rm Sh}$. It shows that the convective-shear waves are excited for $\alpha>1$ at small effective Rayleigh numbers and for $\alpha>1.3$ at large effective Rayleigh numbers. The growth rate of the instability and the frequency of the excited convective-shear waves are very weakly dependent on the effective Rayleigh numbers. Increase of shear, results in increase of the growth rate $\gamma_{\rm inst}$ of the instability and the frequency of the excited convective-shear waves $\omega$. The asymptotic solution of Eq. (13) in the case of $\gamma_{\rm inst}\gg\gamma_{0}$ reads
@@
This instability causes formation of large-scale fluid motions in the form of rolls stretched along the imposed mean wind. This mechanism can also cause the generation of the convective-shear waves with the frequency shown in bottom panels of Figs. 34. The convective-shear waves propagate perpendicular to convective rolls. The predicted motions in convective rolls are characterised by nonzero helicity, in agreement with numerical simulations (see Ref. [28]). Note that similar waves propagating in the direction normal to cloud streets (convective rolls) have been detected in atmospheric convective boundary layers (see Ref. [31]). This large-scale instability are fed by the energy of the convective turbulence.

### v2

@@ Abstract
Based on a mean-field theory of a non-rotating turbulent convection (Phys. Rev. E 66, 066305, 2002), we perform mean-field simulations (MFS) of sheared convection which takes into account an effect of modification of the turbulent heat flux by the non-uniform large-scale motions. This effect is caused by production of additional essentially anisotropic velocity fluctuations generated by tangling of the mean-velocity gradients by small-scale turbulent motions due to the influence of the inertial forces during the lifetime of turbulent eddies. These anisotropic velocity fluctuations contribute to the turbulent heat flux. As the result of this effect, there is an excitation of large-scale convective-shear instability, which causes the formation of large-scale semi-organized structures in the form of rolls. The life-times and spatial scales of these structures are much larger compared to the turbulent scales. By means of MFS performed for stress-free and no-slip vertical boundary conditions, we determine the spatial and temporal characteristics of these structures. Our study demonstrates that the modification of the turbulent heat flux by non-uniform flows leads to a strong reduction of the critical effective Rayleigh number (based on the eddy viscosity and turbulent temperature diffusivity) required for the formation of the large-scale rolls. During the nonlinear stage of the convective-shear instability, there is a transition from the two-layer vertical structure with two rolls in the vertical direction before the system reaches steady-state to the one-layer vertical structure with one roll after the system reaches steady-state. This effect is observed for all effective Rayleigh numbers. We find that inside the convective rolls, the spatial distribution of the mean potential temperature includes regions with a positive vertical gradient of the potential temperature caused by the mean heat flux of the convective rolls. This study might be useful for understanding of the origin of large-scale rolls observed in atmospheric convective boundary layers as well as in numerical simulations and laboratory experiments.
@@ I Introduction
Temperature stratified turbulence and turbulent convection exist in many geophysical and astrophysical flows as well as in industrial flows [see, e.g., Refs. 1, 2, 3, 4, 5, 6, 7]. Turbulence and associated turbulent transport were studied more than 100 years. However, some key questions remain unclear due to extreme values of the governing parameters in geophysical and astrophysical flows [see, e.g., Refs. 8, 9, 10, 11, 12].
@@
  • The level of the mean kinetic energy at saturation increases with increase of the scale separation ratio, and it is very weakly dependent on the effective Rayleigh number.
@@ II Sheared turbulent convection and turbulent heat flux
$\biggl{[}{\partial\over\partial t}+\left(\overline{\bm{U}}+\overline{\bm{U}}_{\rm S}\right)\cdot{\bm{\nabla}}\biggr{]}\overline{\bm{U}}+(\overline{\bm{U}}\cdot{\bm{\nabla}})\overline{\bm{U}}_{\rm S}$
$\quad\quad\quad=-{{\bm{\nabla}}\overline{P}\over\overline{\rho}_{0}}-{\bm{\nabla}}\cdot\hat{\bm{R}}+\beta\,\overline{\Theta}\,{\bm{e}}_{z},$ (1)
$\biggl{[}{\partial\over\partial t}+\left(\overline{\bm{U}}+\overline{\bm{U}}_{\rm S}\right)\cdot{\bm{\nabla}}\biggr{]}\overline{\Theta}=-(\overline{\bm{U}}\cdot{\bm{\nabla}})\overline{T}_{0}-{\bm{\nabla}}\cdot\langle{\bm{u}}\theta\rangle,$
where $\hat{\bm{R}}\equiv R_{ij}=\langle u_{i}\,u_{j}\rangle$ is the Reynolds stress, $\overline{\bm{U}}_{\rm S}=(S\,z,0,0)$ is the shear velocity directed along the $x$ axis, $\overline{\Theta}$ is the mean potential temperature defined as $\overline{\Theta}=\overline{T}\,(\overline{P}_{0}/\overline{P})^{1-1/\gamma}$. Here $\beta=|{\bm{g}}|/T_{0}$ is the buoyancy parameter, ${\bm{g}}=-g\,{\bm{e}}_{z}$ is the acceleration caused by the gravity, ${\bm{e}}_{z}$ is the unit vector in the vertical direction (along the $z$ axis), $\gamma=c_{\rm p}/c_{\rm v}$ is the specific heats ratio, $\overline{T}$ is the mean physical temperature with the reference value $\overline{T}_{0}$ as the temperature in the equilibrium (i.e., the basic reference state), $\overline{P}$ is the mean pressure with the reference value $\overline{P}_{0}$ and $\overline{\rho}_{0}$ is the mean fluid density in the equilibrium. For large Reynolds and Peclet numbers, we neglect in Eqs. (1)–(2) small terms due to the kinematic viscosity and molecular diffusivity of the potential temperature in comparison with those due to the turbulent viscosity and turbulent diffusivity. In Eqs. (1)–(2), the mean fields corresponds to deviations from the equilibrium: $\overline{\bm{U}}_{0}=\overline{\bm{U}}_{\rm S}$, $\;{\bm{\nabla}}\overline{P}_{0}=\overline{\rho}_{0}{\bm{g}}$ and $\overline{\rho}_{0}=$ const.
@@
where ${\bf F}_{x}^{\rm CW}=-\tau_{0}\,({\bf F}_{z}^{\ast}\cdot\mbox{\boldmath$\nabla$})\overline{\bm{U}}_{\rm S}(z)$ is the counter-wind turbulent heat flux and ${\bf F}^{\ast}=-\kappa_{{}_{T}}\mbox{\boldmath$\nabla$}\overline{\Theta}$ is the classical turbulent heat flux, $\tau_{0}$ is the correlation time of turbulent velocity at the integral scale of turbulent motions, $\overline{\bm{W}}=\mbox{\boldmath$\nabla$}{\bf\times}\overline{\bm{U}}$ is the mean vorticity, $\overline{\bm{U}}=\overline{\bm{U}}_{\perp}+\overline{\bm{U}}_{z}$ is the mean velocity with the horizontal $\overline{\bm{U}}_{\perp}$ and vertical $\overline{\bm{U}}_{z}$ components. Equation (3) is derived for large Reynolds and Peclet numbers, which implies that turbulent viscosity and turbulent diffusivity are much larger than the molecular viscosity and molecular diffusivity, respectively. The additional terms in the turbulent heat flux result in the excitation of large-scale instability and formation of the large-scale convective rolls [50, 51, 52].
@@
The physics related to the additional terms in the turbulent heat flux is discussed below. The term $-\tau_{0}\,{\bm{F}}_{z}^{\ast}\,{\rm div}\,\overline{\bm{U}}_{\perp}$ in Eq. (3) for the turbulent heat flux causes the redistribution of the vertical background turbulent heat flux ${\bm{F}}_{z}^{\ast}$ by the perturbations of the convergent (or divergent) horizontal mean velocity $\overline{\bm{U}}_{\perp}$ (see Fig. 1) during the life-time of turbulent eddies. This enhances the vertical turbulent flux of potential temperature due to the converging horizontal motions, which increases the buoyancy, thus creating the upward flow. The latter increases the horizontal convergent flow.
@@
On the other hand, the term $(\tau_{0}/2)\,(\overline{\bm{W}}{\bm{\times}}{\bm{F}}_{z}^{\ast})$ in Eq. (3) produces the horizontal turbulent heat flux by the ”rotation” of the vertical background turbulent heat flux ${\bm{F}}_{z}^{\ast}$ caused by the perturbations of the horizontal mean vorticity $\overline{\bm{W}}_{\perp}$. This decreases local potential temperatures in rising motions, which decreases the buoyancy accelerations, and weakens vertical velocity and vorticity.
@@
The last term $(\tau_{0}/2)\,\left(\overline{\bm{W}}_{z}{\bf\times}{\bf F}_{x}^{\rm CW}\right)$ in Eq. (3) for the turbulent heat flux produces the horizontal heat flux through the ”rotation” of the horizontal background counter-wind turbulent heat flux ${\bf F}_{x}^{\rm CW}$ by the vertical component of the mean vorticity. The counter-wind turbulent flux of potential temperature ${\bf F}_{x}^{\rm CW}$ arises due to the following reasons. In a horizontally homogeneous and sheared convective turbulence, the mean shear velocity $\overline{U}^{\rm S}_{x}(z)$ increases with the height, while the mean potential temperature $\overline{\Theta}(z)$ decreases with the height. Uprising fluid particles produce both, positive fluctuations of potential temperature, $\theta>0$ because $\partial\theta/\partial t\sim-({\bm{u}}\cdot\mbox{\boldmath$\nabla$})\overline{\Theta}(z)$, and negative fluctuations of horizontal velocity, $u_{x}<0$ because $\partial u_{x}/\partial t\sim-({\bm{u}}\cdot\mbox{\boldmath$\nabla$})\overline{U}^{\rm S}_{x}(z)$. This creates a negative horizontal turbulent flux of potential temperature, $u_{x}\,\theta<0$. On the other hand, sinking fluid particles create both, negative fluctuations of potential temperature, $\theta<0$, and positive fluctuations of horizontal velocity, $u_{x}>0$, resulting in negative horizontal turbulent flux of potential temperature, $u_{x}\,\theta<0$. Therefore, the net horizontal turbulent flux of potential temperature is negative, ${\bf F}_{x}^{\rm CW}\equiv\langle u_{x}\,\theta\rangle<0$, in spite of a zero horizontal mean temperature gradient. Therefore, the counter-wind turbulent flux of potential temperature modifies the turbulent potential temperature flux caused by the non-uniform mean velocity field. The counter-wind turbulent flux is associated with non-gradient turbulence transport of heat.
@@
The last term $(\tau_{0}/2)\,\overline{\bm{W}}_{z}{\bf\times}{\bf F}_{x}^{\rm CW}$ in Eq. (3), causes generation of the cross-wind horizontal turbulent heat flux by turning the counter-wind horizontal turbulent flux ${\bf F}_{x}^{\rm CW}$ by perturbations of vertical component of the mean vorticity $\overline{\bm{W}}_{z}$ (see Fig. 2). This produces alternating pairs of convergence or divergence cross-wind turbulent heat fluxes, ${\bf F}_{y}^{\rm new}=(\tau_{0}/2)\,\overline{\bm{W}}_{z}{\bf\times}{\bf F}_{x}^{\rm CW}$, resulting alternative warmer or cooler patches which, in turn, cause alternative upward warm and downward cool motions. This is precisely the mechanism of large-scale instability responsible for formation of the large-scale convective rolls stretched along the mean velocity shear (see Fig. 3) and generation of convective-shear waves propagating perpendicular to the convective rolls in the sheared convection [50, 51, 52].
@@
Figure 3: Large-scale convective rolls stretched along the mean velocity shear and generation of convective-shear waves propagating perpendicular to the convective rolls.
@@ III Governing equations and convective-shear instability
${\partial\tilde{\bm{U}}\over\partial t}+\left[\left(\tilde{\bm{U}}+\tilde{\bm{U}}_{\rm S}\right)\cdot{\bm{\nabla}}\right]\tilde{\bm{U}}+(\tilde{\bm{U}}\cdot{\bm{\nabla}})\tilde{\bm{U}}_{\rm S}$
$\;=-{{\bm{\nabla}}\tilde{P}\over\rho_{0}}+{\rm Ra}_{{}_{T}}\,\tilde{\Theta}\,{\bm{e}}_{z}+\Delta\tilde{\bm{U}},$ (4)
@@
where the non-dimensional mean velocity $\tilde{\bm{U}}$ with ${\rm div}\,\tilde{\bm{U}}=0$, the mean potential temperature $\tilde{\Theta}$ and the mean pressure $\tilde{P}$ are shown with tilde, the flux $\tilde{F}_{z}^{\ast}$ is the nondimensional vertical turbulent background heat flux, the unit vector ${\bm{e}}_{z}$ is directed along vertical $z$ axis, the non-dimensional shear velocity is $\tilde{\bm{U}}_{\rm S}=({\rm Sh}\,\tilde{z},0,0)$, and $\tilde{z}$ is is the non-dimensional vertical coordinate.
@@
where $S=\nabla_{z}\overline{U}^{\rm S}_{x}$ is the linear velocity shear, i.e., $S$ is constant, $u_{c}=(gF_{z}\ell_{0})^{1/3}$ is the convective velocity, $F_{z}$ is the vertical turbulent flux of potential temperature, $\kappa_{{}_{T}}$ is the turbulent (eddy) diffusivity and $\tau_{0}=\ell_{0}/u_{0}$. In Eqs (4)–(5) we have neglected small terms $\sim{\rm O}(\epsilon^{2})$.
@@
Figure 4: The growth rate $\gamma_{\rm inst}$ of the instability (upper panel) and the frequency of the excited convective-shear waves $\omega$ (bottom panel) versus the parameter $\alpha$ for the effective Rayleigh number ${\rm Ra}_{{}_{T}}=0.5$ and different values of the non-dimensional shear number ${\rm Sh}=$ 0.03 (dashed); 0.1 (dashed-dotted); 0.3 (solid).
Figure 5: The growth rate $\gamma_{\rm inst}$ of the instability (upper panel) and the frequency of the excited convective-shear waves $\omega$ (bottom panel) versus the parameter $\alpha$ for the effective Rayleigh number ${\rm Ra}_{{}_{T}}=10^{3}$ and different values of the non-dimensional shear number ${\rm Sh}=$ 0.03 (dashed); 0.1 (dashed-dotted); 0.3 (solid).
@@
Numerical solution of Eq. (13) yields the growth rate $\gamma_{\rm inst}$ of the instability and the frequency $\omega$ of the excited convective-shear waves versus the parameter $\alpha$ for the effective Rayleigh numbers ${\rm Ra}_{{}_{T}}=0.5$ (see Fig. 4) and ${\rm Ra}_{{}_{T}}=10^{3}$ (see Fig. 5), and for different values of the non-dimensional shear number ${\rm Sh}$. It shows that the convective-shear waves are excited for $\alpha>1$ at small effective Rayleigh numbers and for $\alpha>1.3$ at large effective Rayleigh numbers. The growth rate of the instability and the frequency of the excited convective-shear waves are very weakly dependent on the effective Rayleigh numbers. Increase of shear, results in increase of the growth rate $\gamma_{\rm inst}$ of the instability and the frequency of the excited convective-shear waves $\omega$. The asymptotic solution of Eq. (13) in the case of $\gamma_{\rm inst}\gg\gamma_{0}$ reads
@@
This instability causes formation of large-scale fluid motions in the form of rolls stretched along the imposed mean wind. This mechanism can also cause the generation of the convective-shear waves with the frequency shown in bottom panels of Figs. 45. The convective-shear waves propagate perpendicular to convective rolls. The predicted motions in convective rolls are characterised by nonzero helicity, in agreement with numerical simulations (see Ref. [28]). Note that similar waves propagating in the direction normal to cloud streets (convective rolls) have been detected in atmospheric convective boundary layers (see Ref. [31]). This large-scale instability is fed by the energy of the convective turbulence.